OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to -3x + y = -1?

OpenStudy (lgbasallote):

where's c?

OpenStudy (anonymous):

c= -1

Parth (parthkohli):

First take it to the y = mx + b form. \( \color{Black}{\Rightarrow y = 3x - 1}\) Now the slope of a perpendicular line would be the negative reciprocal. I'm not sure what you mean by the c here.

OpenStudy (anonymous):

y=3x+c y=-3x+c y=1/3x+c y=-1/3x+c

OpenStudy (anonymous):

A perpendicular line will always have a slope that is the negative reciprocal of the original line. y = 3x - 1 y = -1/3x - 1 That is a line perpendicular to the line. The "b" in y = mx + b can be changed interchangeably.

OpenStudy (anonymous):

Since in this problem, c is the constant, y = -1/3x + c

OpenStudy (anonymous):

Can you also do the simplified form of sqrt(72x^5y^4)

OpenStudy (anonymous):

or What is the value of x in the solution to the following system of equations? x - 2y = 2 3x + y = 6

OpenStudy (anonymous):

7x = 14 x = 2 √(72x^5y^4) = √(36*2x^5y^4) = 6x^2y^2√2x